OpenStudy (anonymous):

Which of the following exponential functions goes through the points (1, 12) and (2, 36)?

OpenStudy (anonymous):

\[f(x)=4(3)^{x}\] \[f(x)=3(4)^{-x}\] \[f(x)=3(4)^{x}\] \[f(x)=4(3)^{x}\]

OpenStudy (lochana):

just plug in. If the y is the same as what you get from f(x). That is the answer

OpenStudy (lochana):

it is A

OpenStudy (anonymous):

thank you!

OpenStudy (lochana):

which makes D also :)